# Preserving the Hermiticity of the One-Body Density Matrix for a Non-Interacting Fermi Gas

### v2

@@ 2 Hermitian form of the KODM in $d\geq 1$-dimensions
Here $H_{cl}(\mathbf{r},\mathbf{p})=\frac{\mathbf{p}^{2}}{2m}+V(\mathbf{r})$ is the classical counterpart of the quantum one-body Hamiltonian. Substituting equations 2.8, 2.9 and 2.10 into equations 2.5, 2.6 and 2.7 and applying the inverse Laplace transform relationship provided in B by equation 2.4 exactly yield equations 2.2, 2.3 and 2.4. Proof of this is explicitly shown in A.
This offers us the starting point for our symmetrization procedure where we will transform the variables of equations 2.8, 2.9 and 2.10 from $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$ before taking the inverse Laplace transforms in equations 2.5, 2.6 and 2.7. Upon first glance one might attempt to directly transform the KODM [equations 2.1, 2.2, 2.3 and 2.4] into symmetric coordinates, however, by doing this one risks incorrectly neglecting the the appearance of additional terms of the symmetric semiclassical expansion proportional to both $\hbar$ and 2superscriptPlanck-constant-over-$\hbar^{2}$. These new terms form from the $\exp[-\beta V(\mathbf{r})]$ factor within the integrands in equations 2.8, 2.9 and 2.10 which we will show. For simplicity, we will perform the transformation of variables to each term separately, starting with $A_{0}$ and ending with $A_{2}$. The results will be summed afterwards via equation 2.1 to obtain the KODM in terms of the symmetric coordinates $(\mathbf{R},\mathbf{s})$.
2.1 $A_{0}$ change of variables
The $d\geq 1$-dimensional Gaussian integral in equation 2.8 can straightforwardly be evaluated to yield
$C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)=g\bigg{(}\frac{m}{2\pi\beta\hbar^{2}}\bigg{)}^{d/2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}.$ (2.11)
To move into symmetric coordinates we use equation 1.3 to make the substitution $V(\mathbf{r})=V(\mathbf{R}+\mathbf{s}/2)$. Next we Taylor-expand this to second order in $\mathbf{s}$ (equivalently to second order in $\nabla_{\mathbf{R}}$) which yields equation 2.12. Since $\mathbf{s}\rightarrow 0$ is the limit of the local density approximation, and is equivalent to the limit where $\hbar\rightarrow 0$, this approximation is well justified.
@@ 2.1 $A_{0}$ change of variables
Substituting equation 2.12 into equation 2.11, Taylor expanding the exponential with respect to small $\mathbf{s}$, and retaining only second order terms gives
$A_{0}\approx X_{a}+X_{b}+X_{c};$ (2.13)
@@
where the $X_{a,b,c}$ terms have been separated based upon their power of $\beta$. $X_{a}$ originates from the zeroth order part of the exponential expansion, where $\exp[-\beta V(\mathbf{R}+\mathbf{s}/2)]$ is simply replaced by $\exp[-\beta V(\mathbf{R})]$. It is the same in form as equation 2.2 but with $\mathbf{r}$ replaced by $\mathbf{R}$. $X_{b}$, and $X_{c}$ come from the first and second order parts of the exponential expansion, only appearing when the coordinate change is performed before the inverse Laplace transform is taken. Each inverse Laplace transform in equations 2.14, 2.15 and 2.16 can be evaluated via equation 2.4 which is derived in B. After substituting the result and simplifying, the expression for $A_{0}$ in symmetric $(\mathbf{R},\mathbf{s})$ coordinates is obtained and given by
$A_{0}$ $=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z)+\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}$
@@
We also used the definition $k_{F}=k_{F}(\mathbf{R})=\frac{\sqrt{2m}}{\hbar}\big{(}\mu-V(\mathbf{R})\big{)}^{1/2}$ so that $k_{F}$ is now a function of $\mathbf{R}$. This will also be the case for the symmetrized $A_{1}$ and $A_{2}$ terms derived later. The first term in equation 2.17 is precisely in the same form as equation 2.2, as expected, and the three new additional terms are first and second order gradient corrections similar in form to parts of equations 2.3 and 2.4.
2.2 $A_{1}$ change of variables
Before changing the coordinates of equation 2.9, it can be re-expressed by replacing the $\mathbf{p}$ factor with one proportional to $\nabla_{\mathbf{s}}$ and solving the resulting $d\geq 1$-dimensional Gaussian integral in a similar manner to the previous section. This gives
@@ 2.2 $A_{1}$ change of variables
For symmetrization, we again approximate $V(\mathbf{r})$ to second order using equation 2.12. This time we also make use of the approximate gradient identity derived in B, equation 2.8, to transform $\nabla_{\mathbf{r}}V(\mathbf{r})$ into $(\mathbf{R},\mathbf{s})$ coordinates. Combining this with equations 2.18 and 2.6 yields
$A_{1}\approx Y_{a}+Y_{b}+Y_{c};$ (2.19)
@@
Here we have not separated each term based on it’s power of $\beta$ like we did for equations 2.14, 2.15 and 2.16, and instead separate them based on $\nabla_{\mathbf{R}}$. Using the inverse Laplace transformation identity equation 2.4 as before and the gradient and Bessel function relationships of equations 2.7 and 2.3 from B, we can evaluate and simplify equations 2.19, 2.21 and 2.22 to find $A_{1}$ in terms of symmetric coordinates
$A_{1}$ $=-\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}$
$\quad-\frac{gk_{F}^{d-6}z^{4}}{16(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}$
@@
The first term of equation 2.23 is in the same form as equation 2.3 since equation 2.20 comes from the zeroth order part of the expansion of $\exp[-\beta V(\mathbf{r})]$ in equation 2.9, while the additional terms $Y_{b}$ and $Y_{c}$ yield second order gradient corrections. This is similar to the the comments we made regarding equation 2.17.
2.3 $A_{2}$ change of variables
To symmetrizize the final term $A_{2}$, we take a different approach which avoids the need to perform the coordinate transformation before the inverse Laplace transform is evaluated. Since we are only keeping up to second order terms and every term within the integrand of equation 2.10 is already second order in $\nabla_{\mathbf{r}}$, when expanding $\exp[-\beta V(\mathbf{r})]$ using equation 2.12, only the zeroth order part $\exp[-\beta V(\mathbf{R})]$ is retained. Therefore, no additional second order or lower terms appear in the symmetric expression for $A_{2}$. This is in contrast to $A_{0}$ and $A_{1}$ where the additional terms $X_{b,c}$ and $Y_{b,c}$ appeared, respectively. Thus for $A_{2}$, we can simply directly apply our transformation of variables equation 1.3 to the original non-symmetric KODM expression of $A_{2}$, equation 2.4. To do this, we use three gradient identities given by equations 2.8, 2.9 and 2.10 as shown and derived in B. Substituting the identities into equation 2.4 yields the symmetric value of $A_{2}$, which after simplification is given by
$A_{2}$ $\approx\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{R}}^{2}k_{F}^{2}\big{)}+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{R}}k_{F}^{2}\big{)}^{2}$
@@ 2.3 $A_{2}$ change of variables
As expected, this is exactly in the same form as equation 2.4, but with $\mathbf{r}$ replaced by $\mathbf{R}$, similar to how the first terms in equations 2.17 and 2.23 are in the same form as equations 2.2 and 2.3, respectively.
2.4 Symmetrized KODM
We are finally in a position to obtain the symmetric KODM by combining the symmetrized $A_{1,2,3}$ terms. By substituting equations 2.17, 2.23 and 2.24 into equation 2.1, and after simplification, we obtain equation 2.25 which is the key result of this paper. As a reminder, $k_{F}=k_{F}(\mathbf{R})=\frac{\sqrt{2m}}{\hbar}\big{(}\mu-V(\mathbf{R})\big{)}^{1/2}$.
$\rho_{K}(\mathbf{R},\mathbf{s})$ $=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z)+\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{R}}^{2}k_{F}\big{)}$
@@ 2.4 Symmetrized KODM
$\quad+\frac{gk_{F}^{d-4}z^{3}}{48(2\pi z)^{d/2}}J_{\frac{d}{2}-1}\bigg{[}\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}k_{F}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]}.$ (2.25)
@@
Equation 2.25 is exactly identical to the $d\geq 1$-dimensional GVODM originally proposed by reference [1] based upon their $d=1,2,3,4$ derivations. Here we have provided an explicit derivation of their result starting from the known $d\geq 1$-dimensional KODM expression of reference [20], and have shown that upon the proper change to symmetric coordinates, the KODM and GVODM are identical, $\rho_{K}(\mathbf{R},\mathbf{s})=\rho_{GV}(\mathbf{R},\mathbf{s})$. The loss of both Hermiticity and idempotency of the KODM is only an artifact of a non-symmetric truncation of the semiclassical $\hbar$ expansion.
3 Hermitian form of the WKODM in $d\geq 1$-dimensions
3.1 Symmetric Bloch matrix
In references [1, 21, 22], the manifestly Hermitian and idempotent GVODM in symmetric coordinates is derived for $d=1,2,3,4$-dimensions despite the fact that it is through means of a non-Hermitian and non-idempotent WKODM method in terms of non-symmetric coordinates. In this section, we will provide a $d\geq 1$-dimensional derivation completely in terms of symmetric coordinates starting from the same point that reference [10] uses for their non-symmetric WKODM derivation, see their section 4.3 where the non-symmetric Bloch matrix is given by
Note that one can easily derive the presence of the degeneracy factor $g$ by accounting for fermionic spin within the position states and simplifying. $\hat{H}$ is the quantum one-body Hamiltonian, and we also note that the WKODM approximation has yet to be applied. Before changing to the symmetric coordinates, consider the unitary translation operator $\hat{U}(\mathbf{a})=\exp[-(i/\hbar)\hat{\mathbf{P}}\cdot\mathbf{a}]$, where $\hat{\mathbf{P}}=-i\hbar\nabla_{\mathbf{R}}$ is the $\mathbf{R}$-space representation of the momentum operator $\hat{\mathbf{P}}|\mathbf{p}\rangle=\mathbf{p}|\mathbf{p}\rangle$, and $\mathbf{a}$ is an arbitrary position space vector. $\hat{U}(\mathbf{a})$ acts on position space states and scalar functions $f$ as follows [26]
$\hat{U}(\mathbf{a})|\mathbf{R}\rangle$ $=|\mathbf{R}-\mathbf{a}\rangle,$
where in between the fourth and last lines we made the substitution $\hat{H}=\hat{\mathbf{P}}^{2}/(2m)+V(\mathbf{R})$. Making use of the momentum-space completeness relationship $\mathds{1}=\int d^{d}p\;|\mathbf{p}\rangle\langle\mathbf{p}|$ and the inner product relationship $\langle\mathbf{R}|\mathbf{p}\rangle=(2\pi\hbar)^{-d/2}\exp[(i/\hbar)\mathbf{p}\cdot\mathbf{R}]$ allows us to recast equation 3.2 as
At this point, we can perform our symmetric semiclassical WKODM expansion following the same techniques used for the usual non-symmetric WKODM method within reference [10]. The process starts by making the substitution $\langle\mathbf{p}|\textrm{Exp}[-\beta\frac{\hat{\mathbf{P}}^{2}}{2m}-\beta V(\mathbf{R}+\mathbf{s}/2)]|\mathbf{R}\rangle=\exp[-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})]\langle\mathbf{p}|\mathbf{R}\rangle\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$, yielding
where $\tilde{w}$ is a function which obeys
$\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2)-(i/\hbar)\mathbf{p}\cdot\mathbf{R}}\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)=\tilde{u}(\mathbf{R},\mathbf{s},\mathbf{p};\beta),$ (3.5)
and $\tilde{u}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ is a function that obeys the Bloch equation [10]:
$\frac{\partial\tilde{u}}{\partial\beta}+\hat{H}\tilde{u}=0.$ (3.6)
For simplicity, we sometimes suppress the $(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ dependence of both the $\tilde{w}$ and $\tilde{u}$ functions, in a similar form to equation 3.6.
3.2 Symmetric WKODM expansion
Substituting equation 3.5 into equation 3.6, making the second order semiclassical approximation $\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)\approx 1+\hbar\tilde{w}_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)+\hbar^{2}\tilde{w}_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$, and collecting terms based on their order of $\hbar$ yields
$-\frac{\beta^{2}}{2m}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}+\frac{\beta}{2m}\nabla_{\mathbf{R}}^{2}V(\mathbf{R}+\mathbf{s}/2)+\frac{i}{\hbar}(\mathbf{p}\cdot\nabla_{\mathbf{R}})\tilde{w}_{1}+\tilde{w}_{2}\bigg{)}=0.$ (3.7)
$\tilde{w}_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=\bigg{(}\frac{i\beta^{2}}{2m}\bigg{)}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)},$ (3.8)
$\tilde{w}_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=-\frac{\beta^{3}}{4m}\nabla_{\mathbf{R}}^{2}V(\mathbf{R}+\mathbf{s}/2)+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}$
$\quad+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{R}})^{2}V(\mathbf{R}+\mathbf{s}/2)-\frac{\beta^{4}}{8m^{2}}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}.$ (3.9)
Aside from a minus sign difference in equation 3.8, equations 3.8 and 3.9 are the same in form as their non-symmetric counterparts $w_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ and $w_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ respectively, as shown in A by equations 1.4 and 1.5 respectively [10]. To keep our notation consistent with the non-symmetric WKODM method, we utilize the fact that equation 3.4 remains invariant when the integrand undergoes the momentum transformation $\mathbf{p}\rightarrow-\mathbf{p}$ (this transformation does not include changing the sign of the differential element $d^{d}p$). This is shown below by equation 3.10, where $f(\mathbf{p})$ is any function which is either a polynomial or can expressed as a polynomial by a Taylor expansion.
Applying equation 3.10 allows us to re-express equation 3.4 as
where we have defined $w(\mathbf{R},\mathbf{s},\mathbf{p};\beta)=\tilde{w}(\mathbf{R},\mathbf{s},-\mathbf{p};\beta)$ so that $w(\mathbf{R},\mathbf{s},\mathbf{p};\beta)\approx 1+\hbar w_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)+\hbar^{2}w_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ with
$w_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=\tilde{w}_{1}(\mathbf{R},\mathbf{s},-\mathbf{p};\beta)=-\bigg{(}\frac{i\beta^{2}}{2m}\bigg{)}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)},$ (3.12)
$=-\frac{\beta^{3}}{4m}\nabla_{\mathbf{R}}^{2}V(\mathbf{R}+\mathbf{s}/2)+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}$
$\quad+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{R}})^{2}V(\mathbf{R}+\mathbf{s}/2)-\frac{\beta^{4}}{8m^{2}}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}.$ (3.13)
Equations 3.12 and 3.13 are in the same form as the non-symmetric expressions of $w_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ and $w_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$, given by equations 1.4 and 1.5 within A, respectively. Similar to the procedure we followed both there and within Section 2 for the KODM derivation (based on reference [10]), we are finally in a position to derive the symmetric WKODM $\rho_{WK}(\mathbf{R},\mathbf{s})$ by utilizing inverse Laplace transforms and separating terms based on their order of $\hbar$. We find
$\rho_{WK}(\mathbf{R},\mathbf{s})$ $\approx B_{0}+B_{1}+B_{2};$ (3.14)
$B_{0}$ $=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{0}(\mathbf{R},\mathbf{s};\beta)}{\beta}\bigg{\}},$ (3.15)
$B_{1}$ $=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{1}(\mathbf{R},\mathbf{s};\beta)}{\beta}\bigg{\}},$ (3.16)
$B_{2}$ $=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{2}(\mathbf{R},\mathbf{s};\beta)}{\beta}\bigg{\}},$ (3.17)
where each term of the approximate symmetric Bloch matrix $C(\mathbf{R},\mathbf{s};\beta)\approx C_{0}(\mathbf{R},\mathbf{s};\beta)+C_{1}(\mathbf{R},\mathbf{s};\beta)+C_{2}(\mathbf{R},\mathbf{s};\beta)$ is given by equations 3.18, 3.19 and 3.20. These expressions are in the same form as those of the non-symmetric case shown in A by equations 1.1, 1.2 and 1.3.
$C_{0}(\mathbf{R},\mathbf{s};\beta)$ $=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}},$ (3.18)
$C_{1}(\mathbf{R},\mathbf{s};\beta)$ $=-\frac{ig\hbar\beta^{2}}{2m(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta),$ (3.19)
$C_{2}(\mathbf{R},\mathbf{s};\beta)$ $=\frac{g\hbar^{2}}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta).$ (3.20)
From hereon, the calculations and approximations we perform to fully simplify and combine equations 3.12, 3.13, 3.14, 3.15, 3.16, 3.17, 3.18, 3.19 and 3.20 are identical to those used in Section 2 for the KODM method, hence to avoid redundancy, we will not explicitly show each step. After simplification one exactly finds that in symmetric coordinates $B_{0}=A_{0}$ [equation 2.17], $B_{1}=A_{1}$ [equation 2.23], and $B_{2}=A_{2}$ [equation 2.24]. Therefore, combining these results together via equation 3.14, we find that the symmetric WKODM is exactly identical to the symmetric KODM [equation 2.25], which we know is exactly the $d\geq 1$-dimensional GVODM originally proposed in reference [1].

### v3

@@ 2 Hermitian form of the WKODM and KODM in $d$-dimensions
Here $H_{cl}(\mathbf{r},\mathbf{p})=\frac{\mathbf{p}^{2}}{2m}+V(\mathbf{r})$ is the classical counterpart of the quantum one-body Hamiltonian. Substituting equations 2.9, 2.10 and 2.11 into equations 2.6, 2.7 and 2.8 and applying the inverse Laplace transform relationship provided in B by equation 2.4, we prove in A that this exactly yields equations 2.2, 2.3 and 2.4, i.e., we find $A_{0}^{K}=A_{0}^{WK}$, $A_{1}^{K}=A_{1}^{WK}$, and $A_{2}^{K}=A_{2}^{WK}$. It then follows that the ODM of Kirzhnits and Wigner-Kirkwood are found to be identical, $\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})=\rho_{WK}(\mathbf{r},\mathbf{r}^{\prime})$. Notice that this equality holds in $d$-dimensions and in the derivation we used the $d$-dimensional Gaussian integral in equation 2.9 for the momentum $\mathbf{p}$ variable, which straightforwardly yields
$C^{WK}_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)=g\bigg{(}\frac{m}{2\pi\beta\hbar^{2}}\bigg{)}^{d/2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}.$ (2.12)
The above analytical expression is very important since it offers, as we will show, the starting point for our symmetrization procedure. It turns out to be more convenient to first transform the spatial variables in equations 2.9, 2.10 and 2.11 from $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$, and after that perform the inverse Laplace transforms in equations 2.6, 2.7 and 2.8 to finally get the ODM in terms of center of mass and relative variables. Our final goal is to compare the resulting expression denoted now by $\rho_{WK}(\mathbf{R},\mathbf{s})$, with the ODM derived in reference [23] via the Grammaticos-Voros expansion, which is directly expressed in terms of $\mathbf{R}$ and $\mathbf{s}$ variables.
For the sake of pedagogical clarity, we will perform the transformation of variables to each term separately, starting with $A_{0}^{{WK}}$ and ending with $A_{2}^{{WK}}$ (we recall that these terms are identical to the $A_{0,1,2}^{{K}}$ of the Kirzhnitz ODM). The results will be summed afterwards via equation 2.1 to obtain equivalently the WKODM or the KODM in terms of the symmetric coordinates $(\mathbf{R},\mathbf{s})$.
2.1 Change of variables $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$ in $A_{0}^{WK}$
In order to evaluate the inverse Laplace transform of equation 2.6 in symmetric coordinates, we first must transform equation 2.12 via equation 1.3. We consider the Taylor expansion of the potential $V(\mathbf{r})$ around the non-locality $s$, and up to second order derivatives of the potential we have
@@
Now, using the Taylor expansion in equation 2.13, we can write
$\textrm{e}^{-\beta V(\mathbf{r})}\approx\textrm{e}^{-\beta V(\mathbf{R})}\textrm{e}^{-\beta\bigg{(}\frac{1}{2}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}\bigg{)}}.$ (2.14)
Expanding the second exponential in the right side of equation 2.14 and retaining only terms up to the second derivative of the potential $V(\mathbf{R})$ we find
$\textrm{e}^{-\beta V(\mathbf{r})}\approx\textrm{e}^{-\beta V(\mathbf{R})}$ $\Bigg{[}1-\beta\bigg{(}\frac{1}{2}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}\bigg{)}$
$+\frac{\beta^{2}}{8}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}^{2}\Bigg{]}.$ (2.15)
Substituting equation 2.15 into equation 2.12 then allows us to write equation 2.2 as
$A_{0}^{{WK}}\approx X_{a}+X_{b}+X_{c};$ (2.16)
@@
Each inverse Laplace transform in equations 2.17, 2.18 and 2.19 can be evaluated via equation 2.4 which is derived in B. After substituting the result and simplifying, the expression for $A_{0}^{{WK}}$ in symmetric $(\mathbf{R},\mathbf{s})$ coordinates is obtained and given by
$A_{0}^{{WK}}$ $=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z)+\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}$
@@
We also used the definition $k_{F}=k_{F}(\mathbf{R})=\frac{\sqrt{2m}}{\hbar}\big{(}\mu-V(\mathbf{R})\big{)}^{1/2}$ so that $k_{F}$ is now a function of $\mathbf{R}$. This will also be the case for the symmetrized $A_{1}^{{WK}}$ and $A_{2}^{{WK}}$ terms which will be derived later.
2.2 Change of variables $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$ in $A_{1}^{{WK}}$
Before changing the coordinates of equation 2.10, it can be re-expressed using the identity $\nabla_{\mathbf{s}}\textrm{e}^{i\mathbf{p}\cdot\mathbf{s}/\hbar}=\frac{i\mathbf{p}}{\hbar}\textrm{e}^{i\mathbf{p}\cdot\mathbf{s}/\hbar}$, and by solving the resulting $d$-dimensional Gaussian integral in a similar manner to the previous subsection. This gives
@@
For symmetrization, we first make use of the approximate gradient identity derived in B, equation 2.8, to transform $\nabla_{\mathbf{r}}V(\mathbf{r})$ into $(\mathbf{R},\mathbf{s})$ coordinates. Combining this with equations 2.21 and 2.7, while making use of the Taylor expansion equation 2.15 in a similar way to how it was previously used for equation 2.16, yields
$A_{1}^{{WK}}\approx Y_{a}+Y_{b}+Y_{c};$ (2.22)
@@
Here we have not separated each term based on its power of $\beta$ like we did for equations 2.17, 2.18 and 2.19, and instead separate them based on $\nabla_{\mathbf{R}}$. Using the inverse Laplace transformation identity of equation 2.4 as before, along with the gradient and Bessel function relationships of equations 2.7 and 2.3 from B, we can evaluate and simplify equations 2.22, 2.24 and 2.25 to find $A_{1}^{{WK}}$ in terms of symmetric coordinates
$A_{1}^{{WK}}$ $=-\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}-\frac{gk_{F}^{d-6}z^{4}}{16(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}$
@@
2.3 Change of variables $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$ in $A_{2}^{{WK}}$
To symmetrize the final term $A_{2}^{{WK}}$, we take a different approach which avoids the need to perform the coordinate transformation before the inverse Laplace transform is evaluated. Since we are only keeping up to second order terms, and every term within the integrand of equation 2.11 is already second order in $\nabla_{\mathbf{r}}$, when expanding the $\textrm{e}^{-\beta V(\mathbf{r})}$ factor in equation 2.11, as shown by equation 2.15, only the zeroth order part ${\textrm{e}^{-\beta V(\mathbf{R})}}$ is retained. Therefore, no additional second order or lower terms appear in the symmetric expression for $A_{2}^{{WK}}$. This is in contrast to $A_{0}^{{WK}}$ and $A_{1}^{{WK}}$ where the additional terms $X_{b,c}$ and $Y_{b,c}$ appeared, respectively. Thus for $A_{2}^{{WK}}$, we can simply directly apply our transformation of variables equation 1.3 to the original non-symmetric KODM expression of $A_{2}^{{K}}$ (equivalent to the non-symmetric WKODM), equation 2.4. To do this, we use three gradient identities given by equations 2.8, 2.9 and 2.10 as shown and derived in B. Substituting the identities into equation 2.4 yields the symmetric value of $A_{2}^{{WK}}$, which after simplification is given by
$A_{2}^{{WK}}$ $\approx\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{R}}^{2}k_{F}^{2}\big{)}+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{R}}k_{F}^{2}\big{)}^{2}$
@@
2.4 Symmetrized WKODM and KODM
We are finally in a position to obtain both the symmetric WKODM and KODM by combining the symmetrized $A_{0,1,2}^{{WK}}$ terms. Recall that the WKODM is given as $\rho_{WK}(\mathbf{R},\mathbf{s})=A_{0}^{WK}+A_{1}^{WK}+A_{2}^{WK}$, and using the results given respectively in equations 2.20, 2.26 and 2.27 into equation 2.1, after the collection and cancellation of identical terms, we find the final result
${\rho_{WK}(\mathbf{R},\mathbf{s})}=\rho_{K}(\mathbf{R},\mathbf{s})$ $=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z)+\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{R}}^{2}k_{F}\big{)}$
@@ 2.4 Symmetrized WKODM and KODM
$\quad+\frac{gk_{F}^{d-4}z^{3}}{48(2\pi z)^{d/2}}{J_{\frac{d}{2}-1}(z)}\bigg{[}\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}k_{F}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]}.$ (2.28)
@@
The above expression constitutes the key result of this paper, and is exactly identical to the $d$-dimensional GVODM originally proposed by reference [23] based upon their $d=1,2,3,4$ derivations, see equation (4.13) in the aforementioned reference. Here we have provided an explicit derivation of their result starting from the known $d$-dimensional KODM expression of reference [22] in non-symmetric coordinates (equivalent to the WKODM), and have shown that upon the proper change to symmetric coordinates, the WKODM, KODM, and GVODM are identical, ${\rho_{WK}(\mathbf{R},\mathbf{s})=}\rho_{K}(\mathbf{R},\mathbf{s})=\rho_{GV}(\mathbf{R},\mathbf{s})$. The losses of both Hermiticity and idempotency of the Wigner-Kirkwood and Kirzhnits ODMs are only an artifact of non-symmetric truncations of the respective semiclassical $\hbar$-expansions.